OpenStudy (anonymous):

Describe how the graph of y = x^2 can be transformed to the graph of the given equation. y = (x-16)^2

OpenStudy (jdoe0001):

y = A( Bx + C ) + D C = horizontal shift, C > 0, to the left, C < 0, to the right D = vertical shift, D > 0, up, D<0, down

OpenStudy (anonymous):

Shift the graph of y= x^2 right 16 units.

OpenStudy (anonymous):

right side by 16